OpenStudy (erinkb99):

Cleaning pumps remove oil at the rate modeled by the function R, given by R(t)=2+cos(pi*t/11) with t measure in hours and and R(t) measured in gallons per hour. How much oil will the pumping stations remove during the 6-hour period from t = 0 to t = 6?

jimthompson5910 (jim_thompson5910):

So they want you to find the value of \(\LARGE \int_{0}^{6}R(t)dt\)

OpenStudy (erinkb99):

Wait wrong question

jimthompson5910 (jim_thompson5910):

Which means you'll need to find the antiderivative of \(\Large 2+\cos\left(\pi*\frac{t}{11}\right)\)

jimthompson5910 (jim_thompson5910):

you already figured this one out?

OpenStudy (erinkb99):

Yes

OpenStudy (erinkb99):

Let F(x) integral from [0, 2x] tan(t^2) dt. Use your calculator to find F″(1).

jimthompson5910 (jim_thompson5910):

are you familiar with the fundamental theorem of calculus (FTC) ?

OpenStudy (erinkb99):

yes

jimthompson5910 (jim_thompson5910):

If \[\Large f(x) = \int_{a}^{x}g(t)dt\] then \[\Large f \ '(x) = g(x)\]

OpenStudy (erinkb99):

yes, so how does the second derivative relate?

jimthompson5910 (jim_thompson5910):

A slight variation of the rule If \[\Large f(x) = \int_{a}^{h(x)}g(t)dt\] then \[\Large f \ '(x) = g(h(x))*h \ '(x)\] where h(x) is any function of x

jimthompson5910 (jim_thompson5910):

the second derivative is simply the derivative of the first derivative

OpenStudy (erinkb99):

yes, that makes sense

OpenStudy (erinkb99):

so then the derivative of tan(x^2) is 2xsec^2(x^2) right?

jimthompson5910 (jim_thompson5910):

`so then the derivative of tan(x^2) is 2xsec^2(x^2) right?` correct

OpenStudy (erinkb99):

And then just plug in 1?

jimthompson5910 (jim_thompson5910):

Original \[\Large F(x) = \int_{0}^{2x}\tan(t^2)dt\] ------------------------------------------ First Derivative \[\Large F \ '(x) = \tan((2x)^2)*\frac{d}{dx}[2x]\] \[\Large F \ '(x) = \tan(4x^2)*2\] \[\Large F \ '(x) = 2\tan(4x^2)\] ------------------------------------------ Second Derivative \[\Large F \ ''(x) = \frac{d}{dx}\left[F \ '(x)\right]\] \[\Large F \ ''(x) = \frac{d}{dx}\left[2\tan(4x^2)\right]\] \[\Large F \ ''(x) = 2\sec^{2}(4x^2)*\frac{d}{dx}[4x^2]\] \[\Large F \ ''(x) = 2\sec^{2}(4x^2)*8x\] \[\Large F \ ''(x) = 2*8x\sec^{2}(4x^2)\] \[\Large F \ ''(x) = 16x\sec^{2}(4x^2)\]

jimthompson5910 (jim_thompson5910):

what you said is correct, but there were certain pieces missing. What I just posted is how to do the full problem